If P(E) = 0.40, P(EorF) = 0.50 , and P(E and F)=0.10 , find P(F)
Answer by ikleyn(52784)

.

Use the identity


    P(E or F) = P(E) + P(F) - P(E and F).


Substitute the given values


    0.5 = 0.4 + P(F) - 0.1.


From the last equation get


    P(F) = 0.5 - 0.4 + 0.1 = 0.2.      ANSWER

Solved, explained, answered and completed.